# Bipartite entanglement extracted from multimode squeezed light generated in lossy waveguides

### v2

@@ Abstract
Entangled two-mode Gaussian states constitute an important building block for continuous variable quantum computing and communication protocols and are thus of high demand for many experiments. In this work, we study such kind of states which are extracted from multimode light generated via type-II parametric down-conversion (PDC) in lossy waveguides. For such states, we demonstrate that the squeezing quantifies entanglement and we construct a measurement basis which results in the maximal bipartite entanglement. We illustrate our findings by numerically solving the spatial master equation for PDC in a Markovian environment. The optimal measurement modes are compared with two widely-used broadband bases: the Mercer-Wolf basis (the first-order coherence basis) and the Williamson-Euler basis.
@@
Entangled two-mode bipartite Gaussian states play a fundamental role in quantum optics as they represent building blocks for many quantum optical protocols [1, 2]. An efficient source of such highly entangled states is thus an essential part for scalable quantum technological applications [3]. Currently, one of the most promising platforms for producing Gaussian quantum light is based on pulsed parametric down-conversion (PDC) in nonlinear waveguides [4, 5, 6, 7, 8]. However, two main effects are limiting the performance of such sources: scattering of guided light due to the surface roughness [9, 10] and the multimode structure of the generated PDC light [11, 12, 13, 14, 15]. Scattering losses are intrinsic waveguide losses that are caused by technological imperfections during the fabrication. Such type of losses results in the generation of a mixed state in the PDC process. The multimode PDC structure is induced by nonlinear optical coupling of interacting fields and strongly depends on the waveguide dispersion and the pump profile. Two-mode bipartite Gaussian states can be obtained from the initial multimode bipartite light by using a quantum pulse gate (QPG) [16, 17, 18] for each partition. Alternatively, one can selectively work with the chosen two modes of the multimode light by implementing proper gates with the use of homodyne detection [19, 20]. Both of these techniques require the determination of the correct mode shape, since measurements in the wrong modes can lead to the loss of non-classicality [21, 22], which becomes more crucial for PDC light generated in lossy waveguides.
@@
In this paper, we study two-mode bipartite Gaussian states extracted from pulsed type-II multimode PDC generated in lossy waveguides. We show how to choose the measurement basis for the obtaining the maximally-entangled two-mode bipartite state (TMBS) which then can be used for a proper homodyne detection or QPG protocol.
Theoretical model.
@@
For the description of PDC generated in lossy media we use the numerical scheme developed in Refs. [14, 15]. The details of this approach are given in the Appendices A and B. In the framework of multimode Gaussian states, the type-II PDC is described as a multimode bipartite system: part A (also called as signal) and part B (also called as idler), each consisting of $N$ bosonic monochromatic modes, $\hat{\mathbf{a}}=\big{(}\hat{a}_{0},\hat{a}_{1},\dots,\hat{a}_{N}\big{)}^{T}$ and $\hat{\mathbf{b}}=\big{(}\hat{b}_{0},\hat{b}_{1},\dots,\hat{b}_{N}\big{)}^{T}$, respectively. The lower indices correspond to the different frequencies. Having the vacuum state at the input of the waveguide, a solution of the spatial master equations the for correlation matrices for a type-II PDC process in a Markovian environment has the form [15], see also Appendix B,
@@
where the matrices $\braket{\mathbf{\hat{a}^{\dagger}\hat{a}}}$, $\braket{\mathbf{\hat{b}^{\dagger}\hat{b}}}$ and $\braket{\mathbf{\hat{a}\hat{b}}}$ are the correlation matrices with elements $\braket{{\hat{a}^{\dagger}_{i}\hat{a}}_{j}}$, $\braket{{\hat{b}^{\dagger}_{i}\hat{b}_{j}}}$ and $\braket{{\hat{a}_{i}\hat{b}_{j}}}$, respectively. The expectation values $\braket{\mathbf{\hat{a}}^{\dagger}}=\braket{\mathbf{\hat{a}}}=\braket{\mathbf{\hat{b}}^{\dagger}}=\braket{\mathbf{\hat{b}}}\equiv 0$ due to the initial vacuum state. The relationship between the matrices $\mathcal{D}$, $\mathcal{C}$, and the covariance matrix $\Sigma$ are given in Appendix A.
Figure 1: The scheme of a TMBS preparation from a multimode type-II PDC process. After the lossy PDC process, a multimode mixed state with the covariance matrix $\Sigma$ is generated. Two quantum pulse gates (QPG) in partition A with the mode $\mathbf{u}_{A}$ and in partition B with the mode $\mathbf{u}_{B}$, respectively, provide the TMBS with the covariance matrix $\sigma(\mathbf{u}_{A},\mathbf{u}_{B})$.
To extract the maximally-entangled two-mode state from the initial multimode state, we have to find the proper shape of these two broadband modes, which will then be used in the QPGs as shown in Fig. 1 or in homodyne detection for realizing the desired continuous variable protocol. In the general case, such modes represent a superposition of the initial monochromatic PDC modes, therefore, mathematically they can be found using a passive local transformation $U=\big{(}\begin{smallmatrix}\mathbf{u}_{A}&\mathbf{0}_{N}\\ \mathbf{0}_{N}&\mathbf{u}_{B}\end{smallmatrix}\big{)}$, where $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$ are normalized row-vectors of size $N$.
Then, the sought after broadband modes are $\hat{A}=\mathbf{u}_{A}\hat{\mathbf{a}}$ and $\hat{B}=\mathbf{u}_{B}\hat{\mathbf{b}}$. Applying the transformation $U$ to the initial multimode system Eq. (1), via the rules (10), one obtains the following nonzero second-order correlators for the introduced broadband modes:
@@
Note that the modes $\hat{A}$ and $\hat{B}$ can always be chosen to have $\textrm{Im}(\braket{\hat{A}\hat{B}})=0$ via additional phase rotations $\mathbf{u}_{A}\rightarrow\mathbf{u}_{A}e^{i\phi_{A}}$ and $\mathbf{u}_{B}\rightarrow\mathbf{u}_{B}e^{i\phi_{B}}$. The number of photons in the signal and idler parts are $N_{A}=\braket{\hat{A}^{\dagger}\hat{A}}$ and $N_{B}=\braket{\hat{B}^{\dagger}\hat{B}}$, respectively.
By introducing the quadrature operators for broadband modes $\hat{q}^{A}=\hat{A}^{\dagger}+\hat{A}$, $\hat{p}^{A}=i(\hat{A}^{\dagger}-\hat{A})$ and $\hat{q}^{B}=\hat{B}^{\dagger}+\hat{B}$, $\hat{p}^{B}=i(\hat{B}^{\dagger}-\hat{B})$ (with the commutation relations $[\hat{q}^{A},\hat{p}^{A}]=[\hat{q}^{B},\hat{p}^{B}]=2i$, $[\hat{q}^{A},\hat{p}^{B}]=0$), one can construct the covariance matrix $\sigma(\mathbf{u}_{A},\mathbf{u}_{B})$. In what follows, we order the quadratures in this covariance matrix as $(\hat{q}^{A},\hat{p}^{A},\hat{q}^{B},\hat{p}^{B})$ leading to a matrix form, which is a special case of the so-called standard form [23, 24]
@@
where $\alpha\equiv\braket{\hat{p}^{A}\hat{p}^{A}}=\braket{\hat{q}^{A}\hat{q}^{A}}=1+2\braket{\hat{A}^{\dagger}\hat{A}}$, $\beta\equiv\braket{\hat{p}^{B}\hat{p}^{B}}=\braket{\hat{q}^{B}\hat{q}^{B}}=1+2\braket{\hat{B}^{\dagger}\hat{B}}$ and $\gamma\equiv\braket{\hat{q}^{A}\hat{q}^{B}}=-\braket{\hat{p}^{A}\hat{p}^{B}}=2\braket{\hat{A}\hat{B}}$. Note that if $\textrm{Im}(\braket{\hat{A}\hat{B}})=0$, there are no correlations between the coordinates and momenta from different subsystems, namely $\braket{\hat{q}^{A}\hat{p}^{B}}=\braket{\hat{p}^{A}\hat{q}^{B}}=0$, see Appendix A.
To characterize the degree of entanglement of TMBS described by the matrix Eq. (5), we use the logarithmic negativity [25, 26] as the entanglement measure [27, 28]
@@
where the value $\nu_{-}(\tilde{\sigma})$ is the minimal symplectic value of a partially transposed covariance matrix $\tilde{\sigma}=T\sigma T$ with $T=\mathrm{diag}(1,1)\oplus\mathrm{diag}(1,-1)$. Therefore, achieving the TMBS with the maximal degree of entanglement $E_{M}$ is reduced to maximizing the logarithmic negativity by varying the local transformation $U$:
The symplectic spectrum of the partially transformed covariance matrix ${\nu(\tilde{\sigma}})$ can be found as the eigenspectrum of the matrix $|i\Omega\tilde{\sigma}|$ [25, 27, 2], where $\Omega=\omega\oplus\omega$ and $\omega=\big{(}\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\big{)}$. The corresponding characteristic polynomial reads $p(x)=x^{4}-(\alpha^{2}+\beta^{2}+2\gamma^{2})x^{2}+(\alpha\beta-\gamma^{2})^{2}=0$ and results in the symplectic values
@@
On the other hand, the eigenvalues of the covariance matrix $\lambda(\sigma)$ are determined by the roots of its characteristic polynomial, which for the matrix Eq. (5) reads $p(\lambda)=\det(\sigma-\lambda\mathbf{1})=\left((\lambda-\alpha)(\lambda-\beta)-\gamma^{2}\right)^{2}=0$. As can be seen, its roots $\lambda_{\pm}(\sigma)$ are double degenerate and coincide with the symplectic spectrum of the partially transformed covariance matrix $\tilde{\sigma}$; namely, $\lambda_{\pm}(\sigma)\equiv\nu_{\pm}(\tilde{\sigma})$. At the same time, the smallest eigenvalue $\lambda_{-}(\sigma)$ determines the maximal degree of squeezing [29, 30] contained in the system. As long as for the studied system $\lambda_{-}(\sigma)=\nu_{-}(\tilde{\sigma})$, the maximal degree of squeezing is a proper entanglement measure.
This means that the optimization task Eq. (7) can be reformulated in terms of finding the maximal two-mode squeezing of the initial multimode state. Indeed, the lowest eigenvalue $\lambda_{min}$ of the covariance matrix $\Sigma$ determines the maximal squeezing present in the initial multimode system. Then, by determining the vectors $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$, for which $\lambda_{-}(\sigma(\mathbf{u}_{A},\mathbf{u}_{B}))=\lambda_{min}$, one can find the required shapes of the modes that maximize Eq. (7).
In general, the connection between squeezing and entanglement is not as trivial and was studied, e.g., in Ref. [30]: Squeezing is present in entangled Gaussian states, while the opposite is not always correct. However, in our system, there is no single-mode squeezing in either part A or B of the initial state. Indeed, the correlation matrices $\braket{\mathbf{\hat{a}\hat{a}}}=\braket{\mathbf{\hat{b}\hat{b}}}=\mathbf{0}$, see Eq. (1), result in the quadrature variances $\braket{(\Delta\hat{p}^{A})^{2}}\geq 1$ and $\braket{(\Delta\hat{p}^{B})^{2}}\geq 1$. These inequalities hold for any local transformation $U=\big{(}\begin{smallmatrix}\mathbf{u}_{A}&\mathbf{0}_{N}\\ \mathbf{0}_{N}&\mathbf{u}_{B}\end{smallmatrix}\big{)}$, therefore, the only possible squeezing $\braket{(\Delta\hat{p}^{F})^{2}}<1$ can be found only for some joint mode $\hat{F}=t\hat{A}+r\hat{B}$ with $|t|^{2}+|r|^{2}=1$ and $t,r\neq 0$. Such type of squeezing is known as two-mode squeezing. In turn, the mode $\hat{F}$ can always be chosen to get $\braket{(\Delta\hat{p}^{F})^{2}}=\lambda_{-}(\sigma)$.
@@
The basis that provides maximal squeezing of lossy systems and brings the covariance matrix to the second canonical form [29] is the maximally-squeezed basis (MSq-basis) [14]. The first mode $\hat{F}$ of this basis can be found from the eigenvector $\mathbf{v}_{min}=(x^{a}_{1},y^{a}_{1},\dots,x^{a}_{N},y^{a}_{N};x^{b}_{1},y^{b}_{1},\dots,x^{b}_{N},y^{b}_{N})$ that corresponds to the eigenvalue $\lambda_{min}$, namely $\hat{F}=\sum_{n}\big{(}v^{a}_{n}\hat{a}_{n}+v^{b}_{n}\hat{b}_{n}\big{)}$, where we have introduced two complex vectors $[\mathbf{v}^{a}]_{n}=y^{a}_{n}+ix^{a}_{n}$ and $[\mathbf{v}^{b}]_{n}=y^{b}_{n}+ix^{b}_{n}$. Then, the normalized vectors $\mathbf{u}^{MSq}_{A}=\frac{\mathbf{v}^{a}}{|\mathbf{v}^{a}|}$ and $\mathbf{u}^{MSq}_{B}=\frac{\mathbf{v}^{b}}{|\mathbf{v}^{b}|}$ give us the desired broadband modes that minimize Eq. (7). The vectors $\mathbf{u}^{MSq}_{A}$ and $\mathbf{u}^{MSq}_{B}$ define the state called MSq-TMBS and the corresponding covariance matrix $\sigma^{MSq}\equiv\sigma(\mathbf{u}^{MSq}_{A},\mathbf{u}^{MSq}_{B})$. The minimal eigenvalues of matrices $\Sigma$ and $\sigma^{MSq}$ coincide, which guarantees that the generated TMBS is a TMBS with maximal squeezing and, therefore, with maximal entanglement.
Studies of multimode properties of non-classical optical states are commonly performed with two other broadband bases: the Mercer-Wolf (MW) basis [31, 32] (also known as the first-order coherence mode basis [33]) and the Williamson-Euler (WE) basis [2, 34, 35, 36, 37, 38].
Based on the MW and WE decompositions, the vectors $\mathbf{u}^{MW}_{A}$, $\mathbf{u}^{MW}_{B}$ and $\mathbf{u}^{WE}_{A}$, $\mathbf{u}^{WE}_{B}$ can be obtained. These vectors define additional TMBSs (for details see Appendix C) that are further referred to as MW-TMBS and WE-TMBS with the covariance matrices $\sigma^{MW}\equiv\sigma(\mathbf{u}^{MW}_{A},\mathbf{u}^{MW}_{B})$ and $\sigma^{WE}\equiv\sigma(\mathbf{u}^{WE}_{A},\mathbf{u}^{WE}_{B})$, respectively.
@@
Numerical simulations.
Figure 2: Numerical results for a lossless waveguide, the parameters of which are given in the main text. (a) Normalized joint spectral intensity for low-gain PDC, $\delta\omega=\omega-\omega_{0}$, where $\omega_{0}=\omega_{p}/2$ is the central frequency of the PDC light. (b) Dependence of the smallest quadrature variance $\braket{(\Delta\hat{p}^{F})^{2}}=\lambda_{-}(\sigma)$ as a function of the number of photons $N_{A}=N_{B}$. The additional axis shows the corresponding logarithmic negativity $\mathcal{E}(\sigma^{S})$.

### v3

@@ Abstract
Entangled two-mode Gaussian states constitute an important building block for continuous variable quantum computing and communication protocols. In this work, we theoretically study two-mode bipartite states which are extracted from multimode light generated via type-II parametric down-conversion (PDC) in lossy waveguides. For these states, we demonstrate that the squeezing quantifies entanglement and we construct a measurement basis which results in the maximal bipartite entanglement. We illustrate our findings by numerically solving the spatial master equation for PDC in a Markovian environment. The optimal measurement modes are compared with two widely-used broadband bases: the Mercer-Wolf basis (the first-order coherence basis) and the Williamson-Euler basis.
I Introduction
@@
Entangled two-mode bipartite Gaussian states play a fundamental role in quantum optics as they represent building blocks for many quantum optical protocols [1, 2]. An efficient source of highly entangled states is thus an essential part for scalable quantum technological applications [3]. Currently, one of the most promising platforms for producing Gaussian quantum light is based on single-path pulsed parametric down-conversion (PDC) in nonlinear waveguides with second-order nonlinear susceptibility [4, 5, 6, 7, 8]. These platforms potentially allow for the generation of ultrashort pulsed squeezed vacuum states with a high-repetition rate. However, two main factors are limiting the performance of such sources. The first one is the presence of scattering losses, i.e., the intrinsic waveguide losses [9, 10], which are caused by technological imperfections during the fabrication and result in the generation of a mixed state in the PDC process. The second limiting factor is the multimode structure of the generated PDC light [11, 12, 13, 14, 15], which is induced by nonlinear optical coupling of interacting fields and strongly depends on the waveguide dispersion and the pump profile. To accurately operate with multimode PDC light, it is necessary to determine the correct mode shapes, since measurements in the wrong modes can lead to the loss of non-classicality [16, 14, 17]. One of the tasks where mode shapes play a crucial role is the generation of the two-mode bipartite states (TMBS) based on multimode type-II PDC light. Experimentally, TMBS can be obtained from the initial multimode bipartite light by using a single quantum pulse gate (QPG) [18, 19, 20] for each partition. Alternatively, one can selectively work with the chosen two modes of the multimode light by implementing proper gates with the use of homodyne detection [21, 22].
@@
In this paper, we study two-mode bipartite Gaussian states extracted from pulsed type-II multimode PDC generated in lossy waveguides. We show how to choose the measurement basis for obtaining the maximally-entangled two-mode bipartite state (TMBS), which then can be used for a proper homodyne detection or QPG protocol.
II Theoretical description
Theoretical model for type-II PDC.
@@ II Theoretical description
For the description of PDC generated in lossy media, we use the numerical scheme developed in Refs. [14, 15]. The details of this approach are given in the Appendices A and B. In the framework of multimode Gaussian states, the type-II PDC is described as a multimode bipartite system: part A (also called as signal) and part B (also called as idler), each consisting of $N$ bosonic monochromatic modes, $\hat{\mathbf{a}}=\big(\hat{a}_{0},\hat{a}_{1},\dots,\hat{a}_{N}\big)^{T}$ and $\hat{\mathbf{b}}=\big(\hat{b}_{0},\hat{b}_{1},\dots,\hat{b}_{N}\big)^{T}$, respectively. The lower indices correspond to the different discrete frequencies. The quantum state generated via PDC from the vacuum state is fully described by the correlation matrices $\mathcal{D}$ and $\mathcal{C}$, which obey the spatial master equations. For a type-II PDC process in a Markovian environment, see Ref. 15 and Appendix B, their solution has the form
@@
where the matrices $\braket{\mathbf{\hat{a}^{\dagger}\hat{a}}}$, $\braket{\mathbf{\hat{b}^{\dagger}\hat{b}}}$ and $\braket{\mathbf{\hat{a}\hat{b}}}$ are the correlation matrices with elements $\braket{{\hat{a}^{\dagger}_{i}\hat{a}}_{j}}$, $\braket{{\hat{b}^{\dagger}_{i}\hat{b}_{j}}}$ and $\braket{{\hat{a}_{i}\hat{b}_{j}}}$, respectively. The expectation values $\braket{\mathbf{\hat{a}}^{\dagger}}=\braket{\mathbf{\hat{a}}}=\braket{\mathbf{\hat{b}}^{\dagger}}=\braket{\mathbf{\hat{b}}}\equiv 0$ due to the initial vacuum state. Note that the presence of zero matrices $\mathbf{0}$ in Eq. (1) is a property of the considered type-II PDC. Indeed, in this case, the PDC Hamiltonian contains only the interaction terms between the signal and idler subsystems. Below we restrict ourselves to type-II PDC and therefore use Eq. (1) as the starting point for further derivations.
Note that in this study, we only focus on losses during PDC generation; however, the linear external losses, e.g., transmission and detection losses, could easily be included in the consideration. The procedure for introducing frequency-dependent external losses is described in detail in Ref. 14. Here, we only mention that the external losses are included via the transformations $\hat{a}_{n}\rightarrow t^{a}_{n}\hat{a}_{n}$ and $\hat{b}_{n}\rightarrow t^{b}_{n}\hat{b}_{n}$, where $t^{a}_{n}$ and $t^{b}_{n}$ are the complex field transmission coefficients. These transformations do not change the structure of Eq. (1); therefore, all the results we obtained are also valid for frequency-dependent external losses.
Alternatively to the matrices $\mathcal{D}$ and $\mathcal{C}$, the generated PDC state can be presented in terms of the covariance matrix $\Sigma$. The relationship between the matrices $\mathcal{D}$, $\mathcal{C}$, and $\Sigma$ are given in Appendix A.
Figure 1: The scheme of a TMBS preparation from a multimode type-II PDC process. After the lossy PDC process, a multimode mixed state with the covariance matrix $\Sigma$ is generated. Two quantum pulse gates (QPG) in partition A with the mode $\mathbf{u}_{A}$ and in partition B with the mode $\mathbf{u}_{B}$ provide the TMBS with the covariance matrix $\sigma(\mathbf{u}_{A},\mathbf{u}_{B})$.
Two-mode bipartite Gaussian state.
Let us assume that we measure the light in only two modes: one mode in each of subsystems A and B. This scenario can experimentally be realized using the QPGs as shown in Fig. 1 or in homodyne detection, where the modes of local oscillators correspond to measurement modes. As a result, instead of using the initial multimode state, we effectively work with a TMBS built by two measurement modes. In the general case, these modes are represented as a superposition of the initial monochromatic PDC modes, therefore, mathematically they can be found using a passive local transformation
$U=\begin{pmatrix}\mathbf{u}_{A}&\mathbf{0}_{N}\\ \mathbf{0}_{N}&\mathbf{u}_{B}\end{pmatrix}$, (2)
where $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$ are the normalized row-vectors of size $N$. Then the annihilation operators for the broadband measurement modes are given by $\hat{A}=\mathbf{u}_{A}\hat{\mathbf{a}}$ and $\hat{B}=\mathbf{u}_{B}\hat{\mathbf{b}}$.
Applying the transformation $U$ to the initial multimode system Eq. (1), via the rules (17), one obtains the following nonzero second-order correlators for the introduced broadband modes:
@@
To simplify the following expressions, we choose the phases of the modes in such a way that $\textrm{Im}(\braket{\hat{A}\hat{B}})=0$. This can be achieved via additional phase rotations $\mathbf{u}_{A}\rightarrow\mathbf{u}_{A}e^{i\phi_{A}}$ and $\mathbf{u}_{B}\rightarrow\mathbf{u}_{B}e^{i\phi_{B}}$. The number of photons in the signal and idler parts are $N_{A}=\braket{\hat{A}^{\dagger}\hat{A}}$ and $N_{B}=\braket{\hat{B}^{\dagger}\hat{B}}$, respectively. Note that the vectors $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$ fully determine the TMBS: choosing different measurement modes, we obtain different TMBS.
[…]
@@
where we order the quadratures as $(\hat{q}^{A},\hat{p}^{A},\hat{q}^{B},\hat{p}^{B})$. The parameters of the matrix are defined as follows $\alpha\equiv\braket{\hat{p}^{A}\hat{p}^{A}}=\braket{\hat{q}^{A}\hat{q}^{A}}=1+2\braket{\hat{A}^{\dagger}\hat{A}}$, $\beta\equiv\braket{\hat{p}^{B}\hat{p}^{B}}=\braket{\hat{q}^{B}\hat{q}^{B}}=1+2\braket{\hat{B}^{\dagger}\hat{B}}$ and $\gamma\equiv\braket{\hat{q}^{A}\hat{q}^{B}}=-\braket{\hat{p}^{A}\hat{p}^{B}}=2\braket{\hat{A}\hat{B}}$.
Note that the condition $\textrm{Im}(\braket{\hat{A}\hat{B}})=0$ leads to the absence of correlations between the coordinates and momenta of different subsystems, namely $\braket{\hat{q}^{A}\hat{p}^{B}}=\braket{\hat{p}^{A}\hat{q}^{B}}=0$, see Appendix A. Moreover, the covariance matrix Eq. (6) represents a special case of the so-called standard form covariance matrix [23, 24], which significantly simplifies the characterization of entanglement.
Entanglement in TMBS.
Since different broadband modes provide different TMBS, the problem of extracting the TMBS with the maximal degree of entanglement $E_{M}$ reduces to maximizing a proper entanglement measure [25, 26] $\mathcal{E}\left(\sigma\right)$ by varying the local transformation $U$
In this work, as the entanglement measure, we use the logarithmic negativity [27, 28]
@@
[…]
@@
Squeezing in TMBS.
The eigenvalues $\lambda(\sigma)$ of the covariance matrix $\sigma$ are determined by the roots of its characteristic polynomial, which for the matrix Eq. (6) reads $p(\lambda)=\det(\sigma-\lambda\mathbf{1})=\left((\lambda-\alpha)(\lambda-\beta)-\gamma^{2}\right)^{2}=0$. The smallest eigenvalue $\lambda_{-}(\sigma)$ determines the maximal degree of squeezing [29, 30] contained in the system. The states with $\lambda_{-}(\sigma)<1$ are called the squeezed states (in our units, the variance of the vacuum state equals to 1). Note that a higher degree of squeezing corresponds to a lower eigenvalue (lower quadrature variance). In what follows, the squeezing is given in dB as $10\log_{10}(\lambda_{-})$.
Connection between squeezing and entanglement for type-II PDC.
Comparing the eigenvalues of the covariance matrix Eq. (6) with the symplectic spectrum of the partially transformed covariance matrix Eq. (9), we notice that they coincide, namely,
[…]
As a result, the maximal degree of squeezing (minimal eigenvalue $\lambda_{-}(\sigma)$) is a proper entanglement measure for the studied TMBS. This means that the optimization task Eq. (7) can be reformulated in terms of finding the maximal squeezing of the initial multimode state.
Connection between squeezing and entanglement for arbitrary system.
In general, the connection between squeezing and entanglement is not as trivial and was studied, e.g., in Ref. 30: Squeezing is present in entangled Gaussian states, while the opposite is not always true. As an example, we consider the TMBS consisting of two single-mode systems with the covariance matrix $\sigma_{sm}=\big(\begin{smallmatrix}e^{r}&0\\ 0&e^{-r}\end{smallmatrix}\big)\oplus\big(\begin{smallmatrix}e^{r}&0\\ 0&e^{-r}\end{smallmatrix}\big)$. This state is squeezed but is not entangled. Therefore, the presence of single-mode squeezing in one of the subsystems breaks Eq. (10).
To answer the question why Eq. (10) holds for our system, we notice that the initial multimode system Eq. (1) does not contain single-mode squeezing in either part A or B. Indeed, the correlation matrices $\braket{\mathbf{\hat{a}\hat{a}}}=\braket{\mathbf{\hat{b}\hat{b}}}=\mathbf{0}$, which physically comes from the absence of signal-signal (and idler-idler) terms in the type-II PDC interaction Hamiltonian and resulting master equations. As a result, the inequalities $\braket{(\Delta\hat{p}^{A})^{2}}=\braket{(\Delta\hat{q}^{A})^{2}}\geq 1$ and $\braket{(\Delta\hat{p}^{B})^{2}}=\braket{(\Delta\hat{q}^{B})^{2}}\geq 1$ hold for any choice of modes $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$, indicating the absence of single-mode squeezing. Therefore, the only possible squeezing $\braket{(\Delta\hat{p}^{F})^{2}}<1$ can only be present in some joint mode $\hat{F}=t\hat{A}+r\hat{B}$ with $|t|^{2}+|r|^{2}=1$ and $t,r\neq 0$. This type of squeezing is known as two-mode squeezing. In turn, the mode $\hat{F}$ can always be chosen to get $\braket{(\Delta\hat{p}^{F})^{2}}=\lambda_{-}(\sigma)$.
@@
To find the maximally entangled TMBS which can be obtained from the initial multimode system Eq. (1), we have to find the modes with maximal squeezing (minimal eigenvalue) for initial multimode system. Indeed, the lowest eigenvalue $\lambda_{min}$ of the covariance matrix $\Sigma$ determines the maximal squeezing present in the initial multimode system. Then, by determining the vectors $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$, for which $\lambda_{-}(\sigma(\mathbf{u}_{A},\mathbf{u}_{B}))=\lambda_{min}$, one finds the required shapes of the modes that maximize Eq. (7).
The basis that provides maximal squeezing of lossy systems and brings the covariance matrix to the second canonical form [29] is the maximally-squeezed basis (MSq-basis) [14]. The mode $\hat{F}$ of this basis is found from the eigenvector $\mathbf{v}_{min}=(x^{a}_{1},y^{a}_{1},\dots,x^{a}_{N},y^{a}_{N};x^{b}_{1},y^{b}_{1},\dots,x^{b}_{N},y^{b}_{N})$ that corresponds to the eigenvalue $\lambda_{min}$, namely $\hat{F}=\sum_{n}\big(v^{a}_{n}\hat{a}_{n}+v^{b}_{n}\hat{b}_{n}\big)$, where we have introduced two complex vectors $\mathbf{v}^{a}]_{n}=y^{a}_{n}+ix^{a}_{n}$ and $\mathbf{v}^{b}]_{n}=y^{b}_{n}+ix^{b}_{n}$. Then, the normalized vectors $\mathbf{u}^{MSq}_{A}=\frac{\mathbf{v}^{a}}{|\mathbf{v}^{a}|}$ and $\mathbf{u}^{MSq}_{B}=\frac{\mathbf{v}^{b}}{|\mathbf{v}^{b}|}$ give us the desired broadband modes that minimize Eq. (7). The vectors $\mathbf{u}^{MSq}_{A}$ and $\mathbf{u}^{MSq}_{B}$ define the state called MSq-TMBS and the corresponding covariance matrix $\sigma^{MSq}\equiv\sigma(\mathbf{u}^{MSq}_{A},\mathbf{u}^{MSq}_{B})$. The minimal eigenvalues of matrices $\Sigma$ and $\sigma^{MSq}$ coincide, which guarantees that the generated TMBS is a TMBS with the maximal squeezing and, therefore, with the maximal entanglement.
Other broadband bases.
Studies of multimode properties of non-classical optical states are commonly performed with two other broadband bases: the Mercer-Wolf (MW) basis [31, 32] (also known as the first-order coherence mode basis [33]) and the Williamson-Euler (WE) basis [2, 34, 35, 36, 37, 38]. Based on the MW and WE decompositions, the vectors $\mathbf{u}^{MW}_{A}$, $\mathbf{u}^{MW}_{B}$ and $\mathbf{u}^{WE}_{A}$, $\mathbf{u}^{WE}_{B}$ can be obtained. These vectors define additional TMBSs (for details see Appendix C) that are further referred to as MW-TMBS and WE-TMBS with the covariance matrices $\sigma^{MW}\equiv\sigma(\mathbf{u}^{MW}_{A},\mathbf{u}^{MW}_{B})$ and $\sigma^{WE}\equiv\sigma(\mathbf{u}^{WE}_{A},\mathbf{u}^{WE}_{B})$, respectively.
Special case: pure multimode state.
@@
III Numerical simulations
Figure 2: Numerical results for a lossless waveguide WG0, the parameters of which are given in the main text. (a) Normalized joint spectral intensity for low-gain PDC, $\delta\omega=\omega-\omega_{0}$, where $\omega_{0}=\omega_{p}/2$ is the central frequency of the PDC light. (b) Dependence of the smallest quadrature variance $\braket{(\Delta\hat{p}^{F})^{2}}=\lambda_{-}(\sigma)$ as a function of the number of photons $N_{A}=N_{B}$. The additional axis shows the corresponding logarithmic negativity $\mathcal{E}(\sigma^{S})$.